OpenStudy (anonymous):

2 + 5 - - 10 10 is wat

OpenStudy (oaktree):

7/10

OpenStudy (oaktree):

Easy problem. Just add the 2 and the 5.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wat is 5/11 + 4/10

OpenStudy (anonymous):

wat is 5/11 + 4/10 =94/110

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

get the LCD=1ox11=110

OpenStudy (anonymous):

or 5/11 + 4/10 (5/11)(10/10) + (4/10)(11/11) =50/110 + 44/110 =94/110